# On the splitting of weak nearly $\cal C$}-manifolds

### v1

@@ Abstract
Weak metric $f$-manifolds, i.e., the linear complex structure on the contact distribution is replaced by a nonsingular skew-symmetric tensor, defined by V. Rovenski and R. Wolak (2022), allow a new look at the theory of almost contact and metric $f$-manifolds. In this paper, we study the curvature and topology of a structure of this type, called the weak nearly $\cal C$}-structure; specifically, we are guided by the following question: Under what conditions are weak nearly $\cal C$-manifolds locally the Riemannian products? Our theorems generalize results of H. Endo (2005), A. Nicola–G. Dileo–I. Yudin (2018) on nearly cosymplectic manifolds and V. Rovenski (2024) on weak nearly cosymplectic manifolds to the context of weak nearly $\cal C$}-manifolds. Namely, we find conditions under which a $(2n+s)$-dimensional weak nearly $\cal C$}-manifold becomes a Riemannian product, and characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. The consequences of these theorems present new results for nearly $\cal C$}-manifolds.
Keywords: weak nearly $\cal C$}-manifold, weak nearly Kähler manifold, curvature tensor, foliation, Killing vector field.
@@ 1 Introduction
A. Gray introduced in [10] a subclass of almost Hermitian manifolds $(M^{2n},J,\bar{g})$, called nearly Kähler manifolds, using the condition that the symmetric part of $\bar{\nabla}J$ vanishes. D. Blair and D. Showers defined in [5] a subclass of almost contact metric manifolds $(M^{2n+1},{f},\xi,\eta,g)$, called nearly cosymplectic manifolds, using a similar condition that only the symmetric part of $\nabla{f}$ vanishes. These odd-dimensional counterparts of nearly Kähler manifolds are useful in classifying almost contact metric manifolds and have applications in physics [11]. In dimensions greater than 5, a nearly cosymplectic manifold is locally the Riemannian product ${\mathbb{R}}\times\bar{M}^{2n}$ or $B^{5}\times\bar{M}^{2n-4}$, where $\bar{M}$ is a nearly Kähler manifold and $B$ is a nearly cosymplectic manifold, see [14]. Any 5-dimensional nearly cosymplectic manifold has an Einstein metric of positive scalar curvature. For example, the sphere $\mathbb{S}^{5}$ is endowed with a nearly cosymplectic structure induced by the almost Hermitian structure of $\mathbb{S}^{6}$.
The $f$-structure introduced by K. Yano, see [21], on a smooth manifold $M^{2n+s}$ serves as a higher-dimensional analog of almost complex structures ($s=0$) and almost contact structures ($s=1$). The recent interest of mathematicians in the $f$-structure is motivated by the study of the dynamics and integration on contact foliations, $\mathfrak{g}$-foliations, and $s$-cosymplectic manifolds, see [1, 7, 9, 13]. Contact foliations generalize to higher dimensions the flow of the Reeb vector field on contact manifolds, and $\cal K$}-structures are a particular case of uniform $s$-contact structures. An important class of metric $f$-manifolds $(M^{\,2n+s},{f},\vec{\xi},\vec{\eta},g)$, where $\vec{\xi}=(\xi_{1},\ldots,\xi_{s})$ are orthonormal vector fields, $\vec{\eta}=(\eta^{1},\ldots,\eta^{s})$ are dual 1-forms, is given by almost $\cal C$}-manifolds, i.e., $\nabla{f}=0$, see [3]. Any such manifold is locally the product of $\mathbb{R}^{s}$ and a Kähler manifold. Nearly $\cal C$}-manifolds are defined similarly to nearly Kähler manifolds, starting from almost $\cal C$}-manifolds, that is by condition $(\nabla_{X}f)X=0\ (X\in TM)$, e.g., [2]. The question arises: Under what conditions are nearly $\cal C$-manifolds locally the Riemannian products?
@@
In [15, 16], Rovenski et al. introduced metric structures on a smooth manifold that generalize the almost Hermitian structure, almost contact metric structure, and metric $f$-structure. These so-called “weak” metric structures (the linear complex structure on the contact distribution is approximated by a nonsingular skew-symmetric tensor) made it possible to take a new look at the classical structures and find new applications in differential geometry and physics [22]. V. Rovenski defined the weak nearly cosymplectic and weak nearly Kähler structures (generalizing nearly cosymplectic and nearly Kähler structures, respectively), and under certain conditions obtained their decomposition, see [18], generalizing the classical result in [14]. In [20], V. Rovenski initiated the study of a special class of metric $f$-manifolds called weak nearly $\cal C$-manifolds, which generalize nearly $\cal C$-manifolds. Guided by [18, 20], in the paper we investigate the above question for the case of weak nearly $\cal C$-manifolds.
In this article, we study the curvature and topology of weak nearly $\cal C$}-manifolds and find conditions (9) and (13) that are trivially satisfied by almost metric $f$-manifolds and under which weak nearly $\cal C$}-manifolds are locally Riemannian products. In Section 2, following the introductory Section 1, we recall necessary results on weak metric $f$-structures. In Section 3, we formulate auxiliary lemmas on the extrinsic geometry and curvature of weak nearly $\cal C$}- and weak nearly Kähler manifolds.
@@
2 Basic concepts and examples of $f$-structure
In this section, we will review some well-known concepts and results, see [16, 17, 19, 20].
@@ 2 Basic concepts and examples of $f$-structure
A Riemannian manifold $(M,g)$ of even dimension equipped with a skew-symmetric (1,1)-tensor f{f} such that the tensor ${f}^{\,2}$ is negative-definite is called a weak Hermitian manifold. Such (M,f,g)(M,{f},g) is called a weak Kähler manifold if $\nabla{f}=0$, where $\nabla$ is the Levi-Civita connection. A weak Hermitian manifold is called a weak nearly Kähler manifold, if $(\nabla_{X}{f})Y+(\nabla_{Y}{f})X=0\ (X,Y\in\mathfrak{X}_{M})$.
@@
weak metric $f$-structure on a smooth manifold $M^{2n+s}$ $(n,s>0)$ is a set $({f},Q,\vec{\xi},\vec{\eta}$, $g)$, where f{f} is a skew-symmetric $(1,1)$-tensor of rank $2n$, $Q$ is a self-adjoint nonsingular $(1,1)$-tensor, $\vec{\xi}=(\xi_{1},\ldots,\xi_{s})$ are orthonormal vector fields, $\vec{\eta}=(\eta^{1},\ldots,\eta^{s})$ are dual 1-forms, and $g$ is a Riemannian metric on $M$, satisfying
${f}^{2}=-Q+\sum\nolimits_{i=1}^{s}{\eta^{i}}\otimes{\xi_{i}},\quad{\eta^{i}}({\xi_{j}})=\delta^{i}_{j},\quad Q\,{\xi_{i}}={\xi_{i}},$ (1)
$g({f}X,{f}Y)=g(X,Q\,Y)-\sum\nolimits_{i=1}^{s}{\eta^{i}}(X)\,{\eta^{i}}(Y)\quad(X,Y\in\mathfrak{X}_{M}).$ (2)
@@
The exterior derivative of a 1-form $\omega$} is given by
@@
The fundamental $2$-form $\Phi$ on $(M^{2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ is given by $\Phi(X,Y)=g(X,{f}Y)$ for all $X,Y\in\mathfrak{X}_{M}$. Recall the co-boundary formula for exterior derivative $d$ on a $2$-form $\Phi$,
$d\Phi(X,Y,Z)$ $=X\,\Phi(Y,Z)+Y\,\Phi(Z,X)+Z\,\Phi(X,Y)$
$\ -\Phi([X,Y],Z)-\Phi([Z,X],Y)-\Phi([Y,Z],X).$
Note that the above equality yields
@@
A weak metric $f$-structure is called a weak $f$-K-contact structure if $\Phi=d{\eta^{1}}=\ldots=d{\eta^{s}}$ is valid and all characteristic vector fields $\xi_{i}$} are Killing, i.e. $(\pounds_{{\xi_{i}}}\,g)(X,Y)=g(\nabla_{X}{\xi_{i}},Y)+g(\nabla_{Y}{\xi_{i}},X)=0$.
Another important case of a weak metric $f$-structure is a weak almost $\cal K$}-structure, that is $d\Phi=0$. Its special case, is a weak almost $\cal C$}-structure, that is $\Phi$ and $\eta^{i}\ (1\leq i\leq s)$ are closed forms. For $s=1$, weak almost $\cal C$}- manifolds reduce to weak almost cosymplectic manifolds. For weak almost $\cal K$}-manifolds (and their subclass, weak almost $\cal C$}-manifolds), the Reeb distribution $\ker f$ is tangent to a foliation. Moreover, weak almost $\cal C$}-manifolds satisfy the following conditions:
@@
trivial for $s=1$. The following condition is a corollary of (6):
@@
A metric $f$-structure, satisfying (5) and $\nabla f=0$, is a weak almost $\cal C$}-structure. We consider the case when only the symmetric part of $\nabla f$ vanishes.
@@
$(ii)$ To construct a weak metric $f$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$ on the Riemannian product $M=\bar{M}\times\mathbb{R}^{s}$ of a weak almost Hermitian manifold $(\bar{M},\bar{f},\bar{g})$ with $\Omega(X,Y)=\bar{g}(X,\bar{f}Y)$ and a Euclidean space $(\mathbb{R}^{s},dy^{2})$, we take any point $(x,y)$ of $M$ and set
@@
$(iii)$ To construct a weak nearly $\cal C$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$ on the Riemannian product $M=\bar{M}\times\mathbb{R}^{s}$ of a weak nearly Kähler manifold $(\bar{M},\bar{f},\bar{g})$ and a Euclidean space $(\mathbb{R}^{s},dy^{2})$, we take any point $(x,y)$ of $M$ and set, as in Example 1(ii)(ii),
$\xi_{i}=(0,\partial_{\,y^{i}}),\ \ \eta^{i}=(0,dy^{i}),\ \ {f}(X,\partial_{\,y^{i}})=(\bar{f}X,0),\ \ Q(X,\partial_{\,y^{i}})=(-\bar{f}^{\,2}X,\partial_{\,y^{i}}).$
By (5), the Reeb distribution $\ker f$ of weak almost $\cal C$}-manifolds is tangent to a $\mathfrak{g}$-foliation with an abelian Lie algebra. A foliation of dimension $s$ on a smooth connected manifold $M$ is called a $\mathfrak{g}$-foliation [1] (where $\mathfrak{g}$ is a Lie algebra of dimension $s$) if there exist complete vector fields $\xi_{1},\ldots,\xi_{s}$ on $M$ which, when restricted to each leaf, form a parallelism of this submanifold with a Lie algebra isomorphic to $\mathfrak{g}$. Below, we present an interesting way to construct a weak nearly $\cal C$}-structure on a manifold $M$ that has two nearly $\cal C$}-structures with the same Reeb vector fields.
Example 2.
Let a Riemannian manifold $(M^{2n+s},g)$ admit two nearly $\mathcal{C}$}-structures with common Reeb vector fields $\xi_{i}$ and one-forms $\eta^{i}=g(\xi_{i},\cdot)$ for $1\leq i\leq s$. Suppose that $f_{1}\neq f_{2}$ are such that the self-adjoint tensor $\psi:=f_{1}f_{2}+f_{2}f_{1}$ is not identically zero. Then the tensor $f:=(\cos t)\,f_{1}+(\sin t)\,f_{2}$ for small $t>0$ ($t$ is a constant) satisfies the nearly $\mathcal{C}$-condition (8). Indeed,
$(\nabla_{X}f)X$ $=\nabla_{X}(fX)-f(\nabla_{X}X)=\cos t\,(\nabla_{X}f_{1})X+\sin t\,(\nabla_{X}f_{2})X=0.$
Also, observe that $f^{2}=-\mathrm{id}+(\sin t\cos t)\,\psi+\sum\nolimits_{i=1}^{s}\eta^{i}\otimes\xi_{i}$. So, $Q=\mathrm{id}-(\sin t\cos t)\,\psi$, which is positive definite for small $t>0$. Thus, $(f,Q,\vec{\xi},\vec{\eta},g)$ is a weak nearly $\cal C$-structure on $M^{2n+s}$.
The following condition is trivial when $Q=\mathrm{id}$:
$(\nabla_{X}\,Q)Y=0\quad(X\in\mathfrak{X}_{M},\ Y\in{\cal D}).$ (9)
Using (9), we have
$(\nabla_{X}\,Q)Y=\sum\nolimits_{i=1}^{s}{\eta^{i}}(Y)(\nabla_{X}\,Q)\xi_{i}=-\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,\widetilde{Q}\nabla_{X}\,\xi_{i}\quad(X,Y\in\mathfrak{X}_{M}).$ (10)
@@
where $R_{{X},{Y}}=[\nabla_{X},\nabla_{Y}]-\nabla_{[X,Y]}$ is the curvature tensor. We have $g(R_{\,\xi,Z}\,{f}X,{f}Y)=0$ for a nearly cosymplectic manifold $(M^{2n+1},f,\xi,\eta,g)$, see [8]; thus its distribution ${\cal D}=\ker\eta$ is curvature invariant:
$R_{X,Y}Z\in{\cal D}\quad(X,Y,Z\in{\cal D}).$ (12)
For example, the property (12) is true for any distribution $\cal D$} on a real space form. A nearly $\cal C$}-manifold satisfying conditions (5) and (6), also satisfies (12) with ${\cal D}=\bigcap_{\,i=1}^{s}\ker{\eta^{i}}$, see Corollary 1. In our study of weak nearly $\cal C$}-manifolds, we assume a condition weaker than (12):
@@
Obviously, conditions (9) and (13) become trivial for metric $f$-manifolds, i.e., when $Q=\mathrm{id}$.
3 Auxiliary results
In this section, we investigate weak nearly $\mathcal{C}$-manifolds satisfying the conditions (5), (6), (9) and (13), or only some of them. Namely, we prove several lemmas that do not require conditions (9) and (13) in the case of nearly $\mathcal{C}$-manifolds, i.e., $Q={\rm Id}$, and generalize to $s>1$ some well-known properties of weak nearly cosymplectic manifolds ($s=1$).

### v2

@@ Abstract
The interest of mathematicians in metric $f$-manifolds, in particular, almost contact metric manifolds, is motivated by the study of the geometry and dynamics of contact foliations, as well as their applications in physics. Weak metric $f$-manifolds, defined by V. Rovenski and R. Wolak (2022), open a new perspective on classical theory of $f$-manifolds and discover new applications. In this paper, we study manifolds of this type, called weak nearly $\cal C$}-manifolds, which generalize almost $\cal C$}-manifolds. We find conditions under which a $(2n+s)$-dimensional weak nearly $\cal C$}-manifold becomes locally a Riemannian product, and characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. The consequences of these theorems present new results for nearly $\cal C$}-manifolds.
Keywords: weak nearly $\cal C$}-manifold, weak nearly Kähler manifold, curvature tensor, totally geodesic foliation, Killing vector field.
@@ 1 Introduction
Contact geometry has garnered increasing interest of mathematicians due to its significant role in physics, e.g. geometrical optics, mechanics, thermodynamics, geometric quantization, control theory, general relativity, e.g., [13, 14]. An important subclass of almost contact metric manifolds $(M^{2n+1},{f},\xi,\eta,g)$, called nearly cosymplectic manifolds, is defined in [4] by condition that the symmetric part of $\nabla{f}$ vanishes. These odd-dimensional counterparts of nearly Kähler manifolds, see [12], are used in classifying almost contact metric manifolds. Any 5-dimensional nearly cosymplectic manifold has an Einstein metric of positive scalar curvature. For example, the sphere $\mathbb{S}^{5}$ is endowed with a nearly cosymplectic structure induced by the almost Hermitian structure of $\mathbb{S}^{6}$. In dimensions greater than 5, nearly cosymplectic manifolds are locally Riemannian products ${\mathbb{R}}\times\bar{M}^{2n}$ or $B^{5}\times\bar{M}^{2n-4}$, where $\bar{M}$ is a nearly Kähler manifold and $B$ is a nearly cosymplectic manifold, see [16].
The $f$-structure introduced by K. Yano, see [24], on a smooth manifold $M^{2n+s}$ serves as a higher-dimensional analog of almost complex ($s=0$) and almost contact ($s=1$) structures. This structure is defined by a (1,1)-tensor $f$ of rank $2n$ such that $f^{3}+f=0$. The tangent bundle $TM$ splits into two complementary subbundles, $2n$-dimensional $f(TM)$ and $s$-dimensional $\ker f$. To generalize concepts and results from almost contact geometry, geometers have studied various broad classes of metric $f$-manifolds (i.e., the distribution $\ker f$ is parallelizable), e.g., [6, 7, 8]. For example, a metric $f$-manifold is termed an almost $\cal K$}-manifold if $d\Phi=0$, where $\Phi(X,Y):=g(X,{f}Y)$. The distribution $\ker f$ of a $\cal K$}-manifold is tangent to a $\mathfrak{g}$-foliation [1] with flat totally geodesic leaves. An important class of metric $f$-manifolds is given by $\cal C$}-manifolds, hence $\nabla{f}=0$, see [3]. Such manifolds are locally products of $\mathbb{R}^{s}$ and Kähler manifolds. Nearly $\cal C$}-manifolds are defined similarly to nearly Kähler manifolds, starting from $\cal C$}-manifolds, i.e., $(\nabla_{X}f)X=0$, see [2].
The question arises: Under what conditions are nearly $\cal C$-manifolds locally Riemannian products?
Recent interest in the $f$-structure among mathematicians is also motivated by the study of dynamics and integration on contact foliations, $\mathfrak{g}$-foliations, and $s$-cosymplectic manifolds, see [1, 9, 11, 15]. Contact foliations generalize to higher dimensions the flow of the Reeb vector field on contact manifolds, and $\cal K$}-structures are a particular case of uniform $s$-contact structures.
@@
In [18], metric structures on a smooth manifold were introduced that generalize the almost Hermitian structure, the almost contact metric structure, and the metric $f$-structure. These so-called “weak metric structures” (the linear complex structure on the contact distribution is replaced by a nonsingular skew-symmetric tensor) allow us to revisit classical theory and discover new applications in the dynamics and integration on contact foliations, differential geometry of Killing vector fields, totally geodesic foliations and Ricci-type solitons, see [19], and mathematical physics, see [23]. In [20], weak nearly cosymplectic manifolds were defined, and their decomposition was proven under certain conditions, generalizing the classical result of [16]. In  [22] a special class of weak metric $f$-manifolds, called weak nearly $\cal C$-manifolds, which generalize nearly $\cal C$-manifolds, was studied.
@@
Corollary 1.
For a nearly $\cal C$}-manifold the contact distribution is curvature invariant:
$R_{X,Y}Z\in{\cal D}\quad(X,Y,Z\in{\cal D}).$ (1)
The following consequences of Theorems 1 and 2 present new results for nearly $\cal C$}-manifolds.
Corollary 2.
Let a nearly $\cal C$-manifold $M^{\,2n+s}({f},\vec{\xi},\vec{\eta},g)$ satisfy (6) and (7). If the coframe $\vec{\eta}$ is uniform, then $n=2$ and $M$ admits an $f$-K-contact structure $(\hat{f},\vec{\xi},\vec{\eta},g)$ with $\hat{f}=-\nabla\xi_{i}$.
Corollary 3.
Let a nearly $\cal C$}- (non-$\cal C$}-) manifold $(M^{\,2n+s},{f},\vec{\xi},\vec{\eta},g)$ of $n>2$ satisfy (6) and (7). If the coframe $\vec{\eta}$ is uniform, then $M$ is locally isometric (the isometry is global if $M$ is complete simply connected) to one of the Riemannian products
[…]
the induced structure on $\bar{M}$ is nearly Kähler, and the induced structure on $B$ is a nearly $\cal C$}- structure.
2 Basic concepts of weak metric $f$-manifolds and examples
In this section, we will review some well-known concepts and results, see [18, 19, 20, 21, 22].
@@ 2 Basic concepts of weak metric $f$-manifolds and examples
A smooth manifold $M^{2n}$ of even dimension equipped with a Riemannian metric $g$ and a skew-symmetric (1,1)-tensor f{f} of rank $2n$ is called a weak almost Hermitian manifold. Such (M,f,g)(M,{f},g) is a weak Kähler manifold if $\nabla{f}=0$, where $\nabla$ is the Levi-Civita connection. A weak almost Hermitian manifold is called a weak nearly Kähler manifold, if $(\nabla_{X}{f})X=0$ for all $X\in\mathfrak{X}_{M}$.
@@
weak metric $f$-structure on a smooth manifold $M^{2n+s}$ $(n,s>0)$ is a set $({f},Q,\vec{\xi},\vec{\eta}$, $g)$, where $f$ is a skew-symmetric $(1,1)$-tensor of rank $2n$, $Q$ is a self-adjoint nonsingular $(1,1)$-tensor, $\vec{\xi}=(\xi_{1},\ldots,\xi_{s})$ are orthonormal vector fields, $\vec{\eta}=(\eta^{1},\ldots,\eta^{s})$ are dual 1-forms, and $g$ is a Riemannian metric on $M$, satisfying
${f}^{2}=-Q+\sum\nolimits_{i}{\eta^{i}}\otimes{\xi_{i}},\quad{\eta^{i}}({\xi_{j}})=\delta^{i}_{j},\quad Q\,{\xi_{i}}={\xi_{i}},$ (2)
$g({f}X,{f}Y)=g(X,Q\,Y)-\sum\nolimits_{i}{\eta^{i}}(X)\,{\eta^{i}}(Y)\quad(X,Y\in\mathfrak{X}_{M}).$ (3)
@@
The fundamental $2$-form $\Phi$ on $(M^{2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ is given by
$\Phi(X,Y)=g(X,{f}Y)\quad(X,Y\in\mathfrak{X}_{M}).$
Recall the co-boundary formulas for exterior derivative $d$ on a 1-form $\omega$} and a $2$-form $\Phi$,
$d\Phi(X,Y,Z)$ $=X\,\Phi(Y,Z)+Y\,\Phi(Z,X)+Z\,\Phi(X,Y)$
$\ -\Phi([X,Y],Z)-\Phi([Z,X],Y)-\Phi([Y,Z],X),$
@@
Note that the above equality for $d\Phi$ yields
@@
Definition 3.
A weak metric $f$-structure is called a weak $f$-K-contact structure if $\Phi=d{\eta^{1}}=\ldots=d{\eta^{s}}$ is valid and all vector fields $\xi_{i}$} are Killing, i.e., $(\pounds_{{\xi_{i}}}\,g)(X,Y)=g(\nabla_{X}{\xi_{i}},Y)+g(\nabla_{Y}{\xi_{i}},X)=0$. Another important case of a weak metric $f$-structure is a weak almost $\cal K$}-structure, that is $d\Phi=0$. Its special case, is a weak almost $\cal C$}-structure, that is $\Phi$ and $\eta^{i}\ (1\leq i\leq s)$ are closed forms.
For $s=1$, weak almost $\cal C$}-manifolds become weak almost cosymplectic manifolds.
For weak almost $\cal K$}-manifolds (and their subclass, weak almost $\cal C$}-manifolds), the Reeb distribution $\ker f$ is tangent to a foliation. Moreover, weak almost $\cal C$}-manifolds satisfy the following conditions:
@@
trivial for $s=1$. The following condition is a consequence of (7):
@@
Remark 1.
By (6), the distribution $\ker f$ of a weak almost $\cal C$}-manifold is tangent to a $\mathfrak{g}$-foliation with an abelian Lie algebra. A foliation of dimension $s$ on a smooth connected manifold $M$ is called a $\mathfrak{g}$-foliation [1], where $\mathfrak{g}$ is a Lie algebra of dimension $s$, if there exist complete vector fields $\xi_{1},\ldots,\xi_{s}$ on $M$ which, when restricted to each leaf, form a parallelism of this submanifold with a Lie algebra isomorphic to $\mathfrak{g}$.
A metric $f$-structure, satisfying (6) and $\nabla f=0$, is a normal weak almost $\cal C$}-structure, see [21, Theorem 5]. We consider the case when only the symmetric part of $\nabla f$ vanishes.
@@
(ii) To construct a weak nearly $\cal C$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$ on the Riemannian product $M=\bar{M}\times\mathbb{R}^{s}$ of a weak nearly Kähler manifold $(\bar{M},\bar{f},\bar{g})$ with $\Omega(X,Y)=\bar{g}(X,\bar{f}Y)$ and a Euclidean space $(\mathbb{R}^{s},dy^{2})$, we take any point $(x,y)$ of $M$ and set
@@
(iii) Let a Riemannian manifold $(M^{2n+s},g)$ admit two nearly $\mathcal{C}$}-structures with common vector fields $\xi_{i}$ and one-forms $\eta^{i}=g(\xi_{i},\cdot)$ for $1\leq i\leq s$. Let $f_{1}\neq f_{2}$ be such that the self-adjoint tensor $\psi:=f_{1}f_{2}+f_{2}f_{1}$ is not identically zero. Then the tensor $f:=(\cos t)\,f_{1}+(\sin t)\,f_{2}$ for small constant $t>0$ satisfies the condition (9). Observe that $f^{2}=-Q+\sum\nolimits_{i}\eta^{i}\otimes\xi_{i}$ with $Q={\rm Id}-(\sin t\cos t)\,\psi$, which is positive definite for small $t>0$. Thus, $(f,Q,\vec{\xi},\vec{\eta},g)$ is a weak nearly $\cal C$-structure on $M^{2n+s}$.
Let denote the $\cal D$} component of a vector. The following condition is trivial when $Q={\rm Id}$:
$((\nabla_{X}Q)Y)^{\top}=0,\quad(\nabla_{Y}Q)Y=0\quad(X\in TM,\ Y\in{\cal D}).$ (10)
Using the first equality of (10), for any $X,Y\in TM$, we have
$(\nabla_{X}Q)Y$ $=[(\nabla_{X}Q)Y]^{\top}+\sum\nolimits_{i}\eta^{i}((\nabla_{X}Q)Y)\,\xi_{i}$
$=-\sum\nolimits_{i}\{\eta^{i}(Y)\widetilde{Q}\nabla_{X}\xi_{i}+g(\widetilde{Q}\nabla_{X}\xi_{i},Y)\,\xi_{i}\}.$ (11)
Remark 2.
In [19], instead of the overly strict condition (4), the condition (10) should be used.
@@
Proof.
The first part easily follows from [22, Proposition 2]. For the second part, following the proof of [22, Proposition 2], we have $(\mathcal{L}_{\xi_{j}}\,g)(\xi_{k},\cdot)=0$ and
$(\mathcal{L}_{\xi_{j}}\,g)(X,Y)=\eta^{j}\big((\nabla_{X}Q)Y+(\nabla_{Y}Q)X\big)\qquad(X,Y\in{\cal D}).$
By this and the second equality of (10), we obtain $(\mathcal{L}_{\xi_{j}}\,g)(X,Y)=0$ for $X,Y\in{\cal D}$. Combining this with $(\mathcal{L}_{\xi_{j}}\,g)(\xi_{k},\cdot)=0$, we conclude that $\xi_{j}$ is Killing. ∎
@@
where $R_{{X},{Y}}=[\nabla_{X},\nabla_{Y}]-\nabla_{[X,Y]}$ is the curvature tensor. We have $g(R_{\,\xi,Z}\,{f}X,{f}Y)=0$ for a nearly cosymplectic manifold, see [10]; thus its distribution ${\cal D}=\ker\eta$ is curvature invariant, see (1). In our study of weak nearly $\cal C$}-manifolds, we assume a condition weaker than (1):
@@
Obviously, conditions (10) and (13) become trivial for metric $f$-manifolds, i.e., when $Q={\rm Id}$.
3 Main results
Here, we prove a splitting result (Theorem 1) for weak nearly $\cal C$}-manifolds satisfying some additional conditions such as (10) and (13). Theorem 2 characterizes $(4+s)$-dimensional weak nearly $\cal C$}-manifolds.